Shift Summary for Eve shift, 10/13/99

Sydney A. Taegar (taegar@physics.arizona.edu)
Wed, 13 Oct 1999 22:02:00 -0700 (MST)

Shift Summary for Eve shift, 10/13/99

On Shift: Sydney, Wah

Hours of beam delivered at all intensities: 0.75
Hours of beam delivered at nominal intensity: 0.75
Hours of E832 data (special runs not included): 0.75
Hours of special runs (laser scans, muon runs...): 0.

Accounting of lost time:

========================================================================
Runs ended this shift:
Run 14832 - 1.8M events - BEAM_799

Tapes brought to FCC this shift:

UP[A,B,C]70-72,UPS024
UPCR01,UPE004,UPM014,UPX005,UPK008,UPF016-017,UPLA03,UPT005,UPH009

========================================================================
Accesses during shift:

Doug and Wah measured DC Z positions.

========================================================================
Overview of shift:

45 min. of beam followed by Run II TeV studies.

RickT investigated HA scaler rate problems for run range 14700-14823;
details in XXII.091.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

If daqmon hangs again (i.e., spill #/tape fraction stop updating) then

during working hours:
call Margaret at x2625 or Carmenita at x2288

otherwise:
call the help desk at x2345 and have them page the
on-call dart support person